OpenStudy (anonymous):

Which of the following is a soltution for 4m - 5 > -9 or 4m - 5 < 3? m > -9 or m < 3 m > -1 or m < 1/2 All Real Numbers No Solution

OpenStudy (anonymous):

@ajprincess plzz help

OpenStudy (ajprincess):

4m - 5 > -9 Can u solve for m? @charlieberzak

OpenStudy (anonymous):

m being what?

OpenStudy (anonymous):

@ajprincess???

OpenStudy (ajprincess):

do u see a m in the inequalty 4m-5>-9

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

u have to solve for that m.

OpenStudy (anonymous):

yea, but I don't understand how to do that...

OpenStudy (ajprincess):

ok. solving for m means on side of the inequality only m should be there. k?

OpenStudy (anonymous):

ookkaayy

OpenStudy (ajprincess):

nw we have 4m-5>-9 as u can see m is nt alone. to make it alone we have to remove all those numbers joinedd to it. 5 is subtracted from 4m. if we r to remove that 5 we have to perform the opposite operation which is addition. that means u have to add 5 to it to remove -5. k?

OpenStudy (anonymous):

im so of gettin now :)

OpenStudy (anonymous):

sort*

OpenStudy (ajprincess):

so which part didnt u understand?

OpenStudy (anonymous):

i mean, thats all i have to do?

OpenStudy (anonymous):

lol i hate math, it messes with my mind to much

OpenStudy (anonymous):

but wouldn't 4 - 5 = -1 which would lead to -1m?

OpenStudy (anonymous):

@jim_thompson5910 plzz help lol im stuck

OpenStudy (anonymous):

its -1<m<2

jimthompson5910 (jim_thompson5910):

you cannot combine 4m and -5 to get -1 or -1m

OpenStudy (anonymous):

cause interval notation (-1,2)

OpenStudy (anonymous):

oh, but then how do i solve this question?

OpenStudy (anonymous):

i just said it lol

OpenStudy (anonymous):

haha i see but im still a little confused

jimthompson5910 (jim_thompson5910):

you need to isolate m in 4m - 5 > -9

OpenStudy (anonymous):

ok so then what?

jimthompson5910 (jim_thompson5910):

what do you get when you do that

OpenStudy (anonymous):

isolate it? ummm 4 - 5 > -9

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

no, in 4m - 5 > -9, we want to get m all by itself on one side

jimthompson5910 (jim_thompson5910):

so you need to move that -5 over first

jimthompson5910 (jim_thompson5910):

you do this by adding 5 to both sides 4m - 5 > -9 4m - 5+5 > -9+5 4m + 0 > -4 4m > -4 what's next?

OpenStudy (anonymous):

m

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

ok what about m

OpenStudy (anonymous):

well m comes next because 4-4 is 0 so the only thing that is left is m

OpenStudy (anonymous):

i am so lost in this subject :/

jimthompson5910 (jim_thompson5910):

If I told you that 4m = -4, what must be the value of m

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

its well -8?

jimthompson5910 (jim_thompson5910):

4 times SOMETHING = -4... what is that SOMETHING

OpenStudy (anonymous):

4*-8 = -4

jimthompson5910 (jim_thompson5910):

4 times -8 is -4 ?

OpenStudy (anonymous):

that what is says in the calculator....

jimthompson5910 (jim_thompson5910):

time to get a new calculator then... 4 times -8 is -32 I think you meant to say 4 plus -8 = -4

OpenStudy (anonymous):

o right lol

jimthompson5910 (jim_thompson5910):

so what goes in the blank 4*______ = -4

OpenStudy (anonymous):

omg -1???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

i will never get this stupid math

jimthompson5910 (jim_thompson5910):

yes it's -1 since 4 times -1 = -4

jimthompson5910 (jim_thompson5910):

how did you get -1

OpenStudy (anonymous):

i figured it out cause 4*1 equals 4 so it must be the same with the -

jimthompson5910 (jim_thompson5910):

So basically a guess and check? A much quicker way to do it is to divide both sides by 4 4m = -4 4m/4 = -4/4 1m = -1 m = -1

OpenStudy (anonymous):

ohhhh now i see

OpenStudy (anonymous):

so the answer to this question would be m > -9??

jimthompson5910 (jim_thompson5910):

no because 4m = -4 leads to m = -1 4m > -4 leads to m > -1

jimthompson5910 (jim_thompson5910):

you divide both sides by 4 to isolate m to go from 4m > -4 to m > -1 Keep in mind that the sign flips if you're dividing both sides by a negative number (that's not the case here, but keep it in mind)

jimthompson5910 (jim_thompson5910):

so putting everything together, we get this 4m - 5 > -9 4m - 5+5 > -9+5 4m + 0 > -4 4m > -4 4m/4 > -4/4 1m > -1 m > -1

jimthompson5910 (jim_thompson5910):

you need to do the same for the other inequality

jimthompson5910 (jim_thompson5910):

so you need to solve 4m - 5 < 3 for m

OpenStudy (anonymous):

so i + 5 again and add 5 to 3?

OpenStudy (anonymous):

4m - 5+5 < 3 +5

jimthompson5910 (jim_thompson5910):

good, what's next

OpenStudy (anonymous):

change the signs?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

4m + 0 < 8

jimthompson5910 (jim_thompson5910):

good and the +0 goes away leaving you with 4m < 8

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

if 4m < 8 then m < ....

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

i dont really know

jimthompson5910 (jim_thompson5910):

think of it as an equation

jimthompson5910 (jim_thompson5910):

if 4m = 8, then m = ??

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

4 times 4 = 8 ?

OpenStudy (anonymous):

no, i was think + so its 2

jimthompson5910 (jim_thompson5910):

replace m with the value you think it is and see if it's true 4m = 8 4*4 = 8 16 = 8 ...false so m = 4 isn't the solution

OpenStudy (anonymous):

i was thinking + so its 2 cause 4*2 = 8

jimthompson5910 (jim_thompson5910):

yes m = 2 is the solution since 4m = 8 4*2 = 8 ... replace m with 2 (since m = 2) 8 = 8 ... true

jimthompson5910 (jim_thompson5910):

so because m = 2 is a solution to 4m = 8 the solution to 4m < 8 is m < 2 see how it's analogous to solving equations?

OpenStudy (anonymous):

yea :)

jimthompson5910 (jim_thompson5910):

so we found... 4m - 5 > -9 turning into m > -1 and 4m - 5 < 3 turning into m < 2 ---------------- essentially 4m - 5 > -9 or 4m - 5 < 3 turns into m > -1 or m < 2

OpenStudy (anonymous):

wow, ur a great teacher

jimthompson5910 (jim_thompson5910):

thanks, so what did you get for the final answer?

OpenStudy (anonymous):

well, im still trying to figure it out... because the answers are m > -9 or m < 3 m > -1 or m < 1/2 All Real Numbers No Solution

jimthompson5910 (jim_thompson5910):

draw out a number line if you're not sure

jimthompson5910 (jim_thompson5910):

and do the proper graphs/shadings

OpenStudy (anonymous):

i think it b? am i right?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

aww

OpenStudy (anonymous):

i was thinking that

jimthompson5910 (jim_thompson5910):

first off b has m < 1/2, not m < 2 but you can think of m > -1 or m < 2 as something else

jimthompson5910 (jim_thompson5910):

did you try the number line graph

OpenStudy (anonymous):

all real numbers? well because they are -1 and 2?